OpenStudy (rainbow_dash):

Find Sn, for the given arithmetic series. a = –26, d = –9, n = 24

OpenStudy (anonymous):

first find \( a_{24}\) \[\Large a_{24}=a_{1}+23d=-26+(23)(-9)=-233\] now use \[\Large S_{n}=\frac{n}{2}[a_{1}+a_{n}]\] where a1=-26 an=-233 n=24 just plug in the values in the formula given below \[\Large S_{n}=\frac{n}{2}[a_{1}+a_{n}]\]

OpenStudy (rainbow_dash):

ohhhkay thanks i wasn't dividing n by two, but i did everything else. thanks bro!

OpenStudy (anonymous):

you are welcome :)